OpenStudy (abbycross167):

Can someone please help me to answer this math question? Please? I will fan and give a medal... Please help me to simplify this Rational Expression - 3x-2 5x+1 ------ + ----- 6x 9x

OpenStudy (amtran_bus):

You need to find a common denominator. What is the smallest number 6 and 9 have in common? It is 18. Since 6*3=18, you need to multiply the numerator by 3. For the right and side, 9*2 = 18, so you need to multiply the top by 2. \[\frac{ 3(3x-2) }{ 18x }+\frac{ 2(5x+1) }{ 18x }\] Now combine the two fractions to a single one and distribute: \[\frac{ 3(3x-2) + 2(5x+1) }{ 18x }\] \[\frac{ 9x-6 + 10x+2 }{ 18x }\] Now simplify by adding like terms: \[\frac{ 19x-4 }{ 18x }\] Do you understand?